OpenStudy (anonymous):

verify that y=3t+t^2 is a solution of ty'-y=t^2

OpenStudy (amistre64):

what is y'?

OpenStudy (anonymous):

o y prime

OpenStudy (anonymous):

or derivative of y

OpenStudy (amistre64):

you know y, define y' and plug it in

OpenStudy (anonymous):

wow is that it?

OpenStudy (amistre64):

yes, if y is a solution, then when you fill in the parts for y ande y', it should equate to t^2

OpenStudy (amistre64):

if it doesnt, then its not a solution

OpenStudy (amistre64):

prove that: x=3 is a solution to, 3x+2 = 11

OpenStudy (anonymous):

ya i know but y' isnt defined in the problem, 9+2=11

OpenStudy (amistre64):

y is defined, and there is only one way to get y' that i can think of :)

OpenStudy (anonymous):

ya i got it thx, it is a solution

OpenStudy (amistre64):

y=3t+t^2 y' = 3+2t ty'-y=t^2 t(3+2t)-(3t+t^2) = t^2 good, now if they want you to prove it by some other method, fine; but the simplest way to check sis to plug in the information